Curriculum vitae  page 1 of 1

Curriculum vitae

to: The President
fax: +1-202-456-1414

from:??

Date: April 30, 2003

Fax: 1 page

Note: Dear Mr. President, I submit my curriculum vitae as arranged by phone. […]

Yours most sincerely

April 30, 2003  page 1 of 1